Explain why NaOH is not the best test reagent for detecting Pb2+?
Answer #1

Solution:

Pb2+ in the form of lead nitrate (Pb(NO3)2 reacts with NaOH to form white precipitate of lead hydroxide (Pb(OH)2.

Pb(NO3)2 + 2NaOH = Pb(OH)2 (white ppt) + 2NaNO3

Pb(OH)2 is an amphotetic hydroxide, it can react with both acids and bases.

When excess NaOH is added, the white precipitate of Pb(OH)2 is dissolved and forms tertahydroxoplumbate (II).

Pb(OH)2 + 2NaOH = Na2[Pb(OH)4] (aq)

Due to the solubility of lead hydroxide ppt in excess NaOH, it is not considered suitable ion to test in NaOH.